OpenStudy (anonymous):

Stupid decimals to fractions!! -0.38 written as a fraction is -6.59 written as a fraction is 0.135 written as a fraction is -0.8 written as a fraction is 0.6 written as a fraction is 0.9 written as a fraction is 0.87 written as a fraction is

jimthompson5910 (jim_thompson5910):

-0.38 = -38/100 = ??? <<< Reduce this -6.59 = -659/100 = ??? <<< Reduce this 0.135 = 135/1000 = ??? <<< Reduce this Hopefully you're starting to see how I'm getting these

jimthompson5910 (jim_thompson5910):

The rule is: if the last digit of the number is in the tenths place, then the number as a fraction is that number without the decimal over 10 if the last digit of the number is in the hundredths place, then the number as a fraction is that number without the decimal over 100 if the last digit of the number is in the thousandths place, then the number as a fraction is that number without the decimal over 1000 etc etc...

OpenStudy (anonymous):

@jim_thompson5910 How do I reduce it though?

jimthompson5910 (jim_thompson5910):

Are you familiar with finding the GCF?

OpenStudy (anonymous):

Nope

jimthompson5910 (jim_thompson5910):

Let's find the GCF of 38 and 100

jimthompson5910 (jim_thompson5910):

List out the prime factorization of each 38: 2*19 100: 2*2*5*5 Notice the two numbers only have a 2 in common and this is the largest thing they have in common. So the GCF of 38 and 100 is 2

jimthompson5910 (jim_thompson5910):

So we divide both pieces by 2 to get 38/2 = 19 100/2 = 50

jimthompson5910 (jim_thompson5910):

so -38/100 = -19/50 which means -0.38 = -19/50

OpenStudy (anonymous):

Oh, so you take the common number in both then divide to get the out come?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

and you keep going until you can't anymore

OpenStudy (anonymous):

That makes so much sense now. Thank you for that.

jimthompson5910 (jim_thompson5910):

but if you divide out the GCF, then you won't have to worry about dividing further because you will have divided out everything that is common between the two

jimthompson5910 (jim_thompson5910):

you're welcome